OpenStudy (anonymous):

The sum of the digits of a two-digit number is seven. When the digits are reversed, the original number is increased by nine. Find the number

OpenStudy (anonymous):

I Really Need help on these 1

OpenStudy (anonymous):

This is a system of equations problem. Suppose the number is AB. So A + B = 7 10B +A = 10A + B +9 Can you solve that?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

You haven't studied systems of equations?

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

my younger brother is having trouble with this problem and im trying to help him find the answer

OpenStudy (anonymous):

OK, so we'll use one of the methods used for these problems, the substitution method.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Take my first equation and rearrange it to A = 7 -B, then substitute it into the second problem.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

That gives us 10B +7 -B = 10(7 -B) + B +9

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Note that we went from two variables to one. Can you solve that equation?

OpenStudy (anonymous):

No i Cant

OpenStudy (anonymous):

OK, then let's do it...... 10B +7 -B = 10(7 -B) + B +9 9B + 7 = 79 - 9B 18B=72 B = 4 so, (by substitution into the equation A = 7-4 = 3) the number is 34. Check step, reverse the digits, 43 is nine more.

OpenStudy (anonymous):

oh i andurstand now thanks alot that really helped

OpenStudy (anonymous):

No sweat.